# Thermoelectric Enhancement via Electronic and Phononic Channels in Staggered and Non-Staggered Dimerized Quantum Ring

### v1

@@ IV.12 Parameter-space mapping of thermoelectric properties in the staggered potential regime
In Fig. 14, we present the phase diagrams of key thermoelectric (TE) quantities, electrical conductance ($G$), Seebeck coefficient ($S$), electronic thermal conductance ($K_{\text{el}}$), and figure of merit (ZTZT) for the case where a staggered modulation of the on-site potential is included. This analysis aims to uncover how the combined effect of quasiperiodic modulation and disorder strength $W$ influences the TE response.
@@
Panel (c) presents the variation of $K^{el}_{\text{max}}$. Like $G_{\text{max}}$, it exhibits nonmonotonic behavior with increasing $W$, though its suppression at higher $t_{1}/t_{2}$ contrasts with the trend observed in $G_{\text{max}}$. This deviation from the Wiedemann–Franz (WF) law suggests the presence of nontrivial transmission pathways, potentially decoupling heat and charge transport channels favorable for achieving higher ZTZT.
@@ IV.13 Role of contact configuration on the maximization of thermoelectric output
Such interference effects are well-documented in mesoscopic systems and are known to influence the line shape of the transmission function near the Fermi energy. In particular, transmission dips or resonances can arise due to constructive or destructive interference, which can in turn boost the Seebeck coefficient and, consequently, ZTZT.
@@ V Conclusion
In summary, we have conducted a detailed investigation into the thermoelectric behavior of a Su–Schrieffer–Heeger (SSH) ring structure subject to a modulated on-site potential of the Aubry–André–Harper (AAH) type. Both staggered and non-staggered configurations of the disorder were considered, in conjunction with variable hopping parameters. Through systematic exploration of the parameter space, we have identified regimes where thermoelectric performance, particularly the figure of merit ZTZT, is significantly enhanced.
@@
$\bullet$ Enhancement of ZTZT via correlated disorder: The introduction of correlated (AAH-type) disorder modifies the electronic spectrum, introducing asymmetry and energy filtering effects that result in marked improvements in ZTZT. The disorder-induced modulation of the density of states near the Fermi level plays a central role in enhancing thermoelectric efficiency.
@@
$\bullet$ Role of quantum interference in transmission control: The ring geometry inherently supports multiple electronic paths. We explore the effect of drain lead placement, which introduces geometric asymmetry and thereby modifies the interference pattern of electronic wavefunctions. This gives rise to transmission resonances and antiresonances, enhancing ZTZT via constructive interference and energy filtering.
$\bullet$ Optimized ZTZT values under specific conditions: For selective combinations of hopping ratio and disorder strength, we find ZTZT values exceeding 7, indicating excellent thermoelectric performance. These results emphasize the potential of such systems for practical thermoelectric applications at the nanoscale.
@@
$\bullet$ Broader impact and future directions: This work contributes novel insights into the thermoelectric performance of low-dimensional and molecular-scale systems. The demonstrated sensitivity of ZTZT to structural parameters encourages further exploration of engineered disorder, topology, and multi-terminal configurations to unlock enhanced functionality in nanoscale thermoelectrics.

### v2

@@ IV.12 Parameter-space mapping of thermoelectric properties in the staggered potential regime
In Fig. 14, we present the phase diagrams of key thermoelectric (TE) quantities, electrical conductance ($G$), Seebeck coefficient ($S$), electronic thermal conductance ($K_{\text{el}}$), and figure of merit ($ZT$) for the case where a staggered modulation of the on-site potential is included. This analysis aims to uncover how the combined effect of quasiperiodic modulation and disorder strength $W$ influences the TE response.
@@
Panel (c) presents the variation of $K^{el}_{\text{max}}$. Like $G_{\text{max}}$, it exhibits nonmonotonic behavior with increasing $W$, though its suppression at higher $t_{1}/t_{2}$ contrasts with the trend observed in $G_{\text{max}}$. This deviation from the Wiedemann–Franz (WF) law suggests the presence of nontrivial transmission pathways, potentially decoupling heat and charge transport channels favorable for achieving higher $ZT$.
@@ IV.13 Role of contact configuration on the maximization of thermoelectric output
Such interference effects are well-documented in mesoscopic systems and are known to influence the line shape of the transmission function near the Fermi energy. In particular, transmission dips or resonances can arise due to constructive or destructive interference, which can in turn boost the Seebeck coefficient and, consequently, $ZT$.
@@ V Conclusion
In summary, we have conducted a detailed investigation into the thermoelectric behavior of a Su–Schrieffer–Heeger (SSH) ring structure subject to a modulated on-site potential of the Aubry–André–Harper (AAH) type. Both staggered and non-staggered configurations of the disorder were considered, in conjunction with variable hopping parameters. Through systematic exploration of the parameter space, we have identified regimes where thermoelectric performance, particularly the figure of merit $ZT$, is significantly enhanced.
@@
$\bullet$ Enhancement of $ZT$ via correlated disorder: The introduction of correlated (AAH-type) disorder modifies the electronic spectrum, introducing asymmetry and energy filtering effects that result in marked improvements in $ZT$. The disorder-induced modulation of the density of states near the Fermi level plays a central role in enhancing thermoelectric efficiency.
@@
$\bullet$ Role of quantum interference in transmission control: The ring geometry inherently supports multiple electronic paths. We explore the effect of drain lead placement, which introduces geometric asymmetry and thereby modifies the interference pattern of electronic wavefunctions. This gives rise to transmission resonances and antiresonances, enhancing $ZT$ via constructive interference and energy filtering.
$\bullet$ Optimized $ZT$ values under specific conditions: For selective combinations of hopping ratio and disorder strength, we find $ZT$ values exceeding 7, indicating excellent thermoelectric performance. These results emphasize the potential of such systems for practical thermoelectric applications at the nanoscale.
@@
$\bullet$ Broader impact and future directions: This work contributes novel insights into the thermoelectric performance of low-dimensional and molecular-scale systems. The demonstrated sensitivity of $ZT$ to structural parameters encourages further exploration of engineered disorder, topology, and multi-terminal configurations to unlock enhanced functionality in nanoscale thermoelectrics.